OpenStudy (sleepyjess):

Given the following reaction: 4Fe+3O2→2Fe2O3 How many grams of Fe2O3 will be produced from 34g of Fe and 12g of O2?

OpenStudy (zpupster):

(\[(34g of Fe)(\frac{ 1 mol Fe}{ 55.847 g Fe })(\frac{ 2 mol Fe2O3 }{ 4 mol Fe })(\frac{ 159.7 g Fe2O3 }{ 1mol })=?\]